A bat can detect small objects such as an
insect whose size is approximately equal to
one wavelength of the sound the bat makes.
If bats emit a chirp at a frequency of
54.5 kHz, and the speed of sound in air is
339 m/s, what is the smallest insect a bat can
detect?
Answer #1
As bat can detect small objects such as an insect whose size is approximately equal to the wavelength of the sound the bat makes,

the smallest insect a bat can detect will be of size approximately equal to the wavelength of the sound the bat make.

The frequency of the sound produced f = 54.5 kHz,

The speed of the sound in air is v = 339 m/s

From the relation

v = f?
The wavelength of the sound will be
? = v/f
? = 339/54.5*10^3
? = 6.22*10^-3 m
? = 6.22 mm
the smallest insect a bat can detect will be of size is 6.22 mm